OpenStudy (anonymous):

Choose the correct simplification of a^5b^6/a^4b^4 A. a^9b^10 B. ab^2 C. 1/ab^2 D. 1/a^9b^10

OpenStudy (anonymous):

B

OpenStudy (anonymous):

B is the correct answer. How to find it is easy! You simply remember the rule to minus exponents with like terms when dividing, and add them when multiplying. Ex) a^5/a^2= a^3, (b^2)(b^7)= b^9